OpenStudy (anonymous):

how did he solve the problem y = (x+1)/(x-2) to get x= 2y+1/y-1?

OpenStudy (anonymous):

make x subject of the formula

OpenStudy (anonymous):

x+1=y(x-2) x+1=yx-2y x-yx=-2y-1 x(1-y)=-1(2y+1) next u proceed

OpenStudy (anonymous):

y(x-2) = x+1 yx-2y =x+1 yx-x = 1+2y x(y-1) =1+2y x=??

OpenStudy (anonymous):

-x=(2y+1)/(1-y)

OpenStudy (anonymous):

@paris818 stop using negatives when you don't have to,they'll confuse you further. look at my work above

OpenStudy (anonymous):

Okay, I was looking @ sail 234

OpenStudy (anonymous):

Thank you.